Write the vector v=( 1, -2, 5) as a linear combination of the vector e1=( 1, 1, 1), e2= ( 1, 2, 3) and e3=( 2, -1, 1)
1
Expert's answer
2020-11-10T06:59:59-0500

Let us assume "v=c_1e_1 + c_2e_2 + c_3e_3." We may write a linear system for every coordinate:

"\begin{cases}
1 = c_1+c_2+2c_3, \\
-2 = c_1+2c_2-c_3, \\
5=c_1+3c_2+c_3.
\end{cases}"

We should solve this linear system and get "c_1 = -6, \; c_2 = 3, \; c_3 =2."

Therefore, "v = -6e_1+3e_2+2e_3."